Write an equation in standard form of the circle described.

Ends of diameter at (−4, −4) and (6, 8)
Answer #1

let olhik) be Al-4,-4) Olhik) 7 B(6,8) unter of cirere h = -4+6 2 - 2 K = – 4 + 8 2 Center, (hik) = (1,2) Rodius, &= √(-4-1)